# Industrial Internet of Things Intelligence Empowering Smart Manufacturing: A Literature Review

### 02 Dec 2023 v1

@@ Abstract
The fiercely competitive business environment and increasingly personalized customization needs are driving the digital transformation and upgrading of the manufacturing industry. IIoT intelligence, which can provide innovative and efficient solutions for various aspects of the manufacturing value chain, illuminates the path of transformation for the manufacturing industry. It’s time to provide a systematic vision of IIoT intelligence. However, existing surveys often focus on specific areas of IIoT intelligence, leading researchers and readers to have biases in their understanding of IIoT intelligence, that is, believing that research in one direction is the most important for the development of IIoT intelligence, while ignoring contributions from other directions. Therefore, this paper provides a comprehensive overview of IIoT intelligence. We first conduct an in-depth analysis of the inevitability of manufacturing transformation and study the successful experiences from the practices of Chinese enterprises. Then we give our definition of IIoT intelligence and demonstrate the value of IIoT intelligence for industries in fucntions, operations, deployments, and application. Afterwards, we propose a hierarchical development architecture for IIoT intelligence, which consists of five layers. The practical values of technical upgrades at each layer are illustrated by a close look on lighthouse factories. Following that, we identify seven kinds of technologies that accelerate the transformation of manufacturing, and clarify their contributions. Finally, we explore the open challenges and development trends from four aspects to inspire future researches.
@@ I Introduction
Therefore, in this paper, we provide a comprehensive overview of IIoT intelligence. Considering that the development of IIoT intelligence integrates the wisdom of governments, enterprises, and researchers, we expand the information sources from academic papers to government documents, industry reports, and other related literature. The essential of this paper is to answer four questions as follows.
@@
To answer the second question, we propose a hierarchical development architecture to build a system-level understanding about how IIoT intelligence empowers smart manufacturing. Specifically, the architecture consists of five layers. Each layer has own missions. The equipment layer builds the basis of automatic production. The network layer connects human-machine-things, and overcomes isolated islands of information. The software layer provides digital representation of industrial processes. The modeling layer makes digital twin for physical processes to connect virtual and physical spaces. The analysis and optimization layer is responsible for mining industrial big data and optimizing industrial processes. Each layer of IIoT intelligence has its own key technologies, making its capabilities stronger. Interactions will occur between levels to fully enhance the intelligence and digitization level of the manufacturing industry. In addition, we also closely observe the lighthouse factory [28][29] to demonstrate that technological upgrades at each layer enhance the company’s returns.
@@
This paper is structured as follows. Section II presents background and transformation secrets of smart manufacturing. The connotation of IIoT intelligence is explained in Section III. Section IV provides a reference hierarchical development architecture to build a systematic understanding for IIoT intelligence. Section V demonstrates the significant effects of IIoT intelligence through case studies on lighthouses and gradually released collaborations. Section VI identities some of the most influential technologies that facilitate IIoT intelligence. Section VII summarizes open challenges and future trends of IIoT intelligence. The conclusion is given in Section VIII.
@@ II-A Background
Recent years have witnessed the rapid development of a large number of advanced intelligent technologies, including cloud computing, 5G connection, industrial data analysis, industrial internet of things, etc. These technologies light the way for upgrading manufacturing. Many countries around the world also support intelligent transformation of manufacturing in the perspectives of policies. Around 20 economies released approximately 100 strategic planning or policy documents on artificial intelligence in the past three years by 2019, among which more than 50 documents are related with industrial application. For example, the governments of the United States, Japan, Germany, China have respectively launched The National Artificial Intelligence Research and Development strategic Plan [2], New Robot Strategy [3], Industrial Strategy 2030: Guidelines for a German and European industrial policy [4], Made in China 2025 [5]. In the government plans, some research domains are highlighted, such as product life cycle optimization, advanced robot, automatic driving, big data mining, cloud computing and other industrial applications.
@@ II-B Secrets to smart manufacturing
In this section, we will answer the question, how to accelerate the progress of intelligent transformation for manufacturing industry, based on the practical experiences of Foxconn Industrial Internet, which is a famous enterprise that has realized a high-level intelligent transformation in China.
Foxconn Industrial Internet summarizes its recipes into three pillars and two capabilities [31], as presented in Fig. 3.
@@
Actually, manufacturing in some countries, such as the United States, Germany, Japan and South Korea, is also facing similar purgatories as the manufacturing in China. Therefore, the recipe of Foxconn Industrial Internet has the potential to be scaled to global manufacturing companies.
@@ III-B How does IIoT intelligence achieve impacts in manufacturing ?
IIoT intelligence can provide some innovative solutions for segments of the value chain of manufacturing, and support to build the pillars and capabilities (shown in Fig. 3). The solutions are able to improve direct labour productivity and employee engagement, raise product quality and overall equipment effectiveness (OEE), reduce R&D costs and energy consumption, etc. We present a close look at two cases to demonstrate the idea as follows.
@@
III-B1 Digital experiments to reduce product R&D costs
Traditionally, the R&D of industrial products should experience five steps: project approval, product research, production preparation, process optimization, and declaration of announcement catalogue. At the stage of product approval, the product development mission statement and R&D budget are required, and the documents need the approval of the president. While researching and developing new products, design drawings and documents of products are required, and trial production of products with some success is expected as well. The production preparation aiming at putting forward production technologies and process planning, and piloting the solutions, according to the design drawings and documents of products. Missions of the process optimization are improving initial process planning, organizing production, and discussing employee income. In the phase of declaration of announcement catalogue, the product declaration planning should be determined according to the product development specification and sales plan.
@@ III-B1 Digital experiments to reduce product R&D costs
Revising the processes of the R&D of products, we find several rounds of trial production, testing and optimization in the physical production lines are important for achieving a new product. However, each round of the physical experiment costs a lot of manpower, material, money and time. Therefore, the traditional R&D process is time-consuming, expensive and environmentally unfriendly.
IIoT intelligence can reduce costs of product R&D by transferring part of the physical experiments into digital world. Specifically, IIoT intelligence is able to take full advantage of knowledge from historical experience and data, then map industrial processes into the digital world, where some trial and error experiments can be carried out at low costs. Besides that, some faults are possible to be predicted automatically according to industrial knowledge, mechanisms and historical experience in the digital world. Some improvement suggestions are likely to be given by some recommendation technologies. Receiving the detected/predicted faults and some valuable ideas, the project team is able to immediately revise the designed products and optimize the architecture of new products.
@@
III-B2 Remote monitoring and control to increase labour efficiency
Traditionally, workers are expected to physically participate in production or monitor equipment operation in the workshop, in order to detect and correct errors during production in time. The maintenance of production equipment depends on the on-site service of technicians.  The deployment of production lines also relies on manual analysis and expert experience.
IIoT intelligent connection and perception technologies can connect equipment, workers and intermediate products in the workshops. It enables remote monitoring of production conditions and remote control of manufacturing equipment, which improve safety through unmanned operation and reduce the number of workers on dangerous sites. The technology can also help improve OEE (overall equipment effectiveness) and labour efficiency by boosting one-man-one-control to one-man-multi-control. With the help of remote monitoring and control, managers are able to flexibly schedule workers according to the progress on-site. In addition, the technologies introduce some white-collar jobs, allowing workers can work in a comfortable office. In this way, employment willingness to the manufacturing industry can be improved and the problem of labour shortage in the manufacturing industry can be alleviated.
@@ III-C The overall effect of IIoT intelligence on smart manufacturing
  • Functions: IIoT intelligence can help companies connect human-cyber-physical spaces, collect industrial data synchronously and accurately, build effective data analysis models, propose scenario-oriented optimization methods, develop microservice architecture, etc.
  • Operation: IIoT intelligence builds cloud manufacturing business model for enterprises to enhance returns [33, 34, 35]. Cloud manufacturing create intelligent factory networks that encourage effective collaboration, through encouraging enterprises to encapsulate distributed manufacturing resources into cloud services and managing them in a centralized way. Clients can use and pay for the cloud manufacturing resources according to their requirements.
  • Deployment: IIoT intelligence provides the cloud-edge-terminal architecture to deploy computing resources in a cost-friendly way. Except that, IIoT intelligence can give efficient production lines deployment suggestions to improve factory output.
  • Application scenarios: IIoT intelligence can be used to solve specific problems in some application scenarios, such as predictive and remote maintenance, product and equipment monitoring, product design feedback optimization, product process traceability, renting manufacturing capacity, product quality inspection, etc.
@@ IV Hierarchical development architecture of IIoT intelligence
To build a systematic understanding of how IIoT intelligence deals with issues in the transformation process of smart manufacturing, this paper presents a hierarchical development architecture for IIoT intelligence. The architecture consists of equipment, network, software, model and optimization from bottom to top, as shown in Fig. 4. The classic architecture of IIoT usually include the bottom two layers [8, 9, 11], however, it cannot bear the weight of the various intelligence that have emerged in recent years. Therefore, we extends software layer, modeling layer, and analysis and optimization layer over the classic architecture.
The functions of each layer in the hierarchical architecture are as follows:
@@ IV-A1 Motivations
Facing the crisis of labour shortage and high labour cost, the manufacturing industry’s desire for a machine that can work like or even better than humans has become increasingly strong, thus promoting the invention and innovation of industrial robots. Except that, there are other reasons that motivate enterprises [36]. a) There are many dull, dirty, dangerous and/or delicate tasks that can be done by robots, which can improve health, safety and job satisfaction of employees. b) The rapid development of robotics technology is continuously improving the automated and flexible production capacity of robots, while also reducing the procurement and maintenance costs. c) Modern robots will support a smaller carbon footprint [37][38], which is profitable for building low-carbon factories according to national requirements. d) More and more robots support plug-and-play integration and their control programming becomes much easy, which allow operators to become proficient in using robots with general training [39]. e) Human-robot collaboration will strengthen flexible production [40][41]. In addition, the aging society of the world has also released the demand for liberation from manual labor. Under the circumstances, governments also propose some important policies111details of the policies are summarized in https://ifr.org/ifr-press-releases/news/investment-in-robotics-research-global-report-2021 to promote development of industrial robots, such as Made in China 2025, the New Robot Strategy In Japan, The Intelligent Robot Development and Supply Promotion Act of Korea, The new European Framework Program Horizon Europe,Germany’s High-Tech Strategy 2025, The National Robotics Initiative (NRI) in the USA.
Except robots, in order to achieve precise perception of the state of the production environment and high-precision control of the production process, it is necessary to introduce smart sensors to improve five senses222taste, touch, smell, sight, hearing of industry. In addition, with the deepening of industrial digital transformation, a large number of industrial data storage and analysis tasks have emerged. To complete these tasks, massive computing and storage resources are required.
@@ IV-A2 Connotation
An industrial robot refers to a robot system that can be used for manufacturing. Industrial robots are automated, programmable and capable of movement on three or more axes. Industrial robots are gradually replacing human to overtake dangerous and delicate tasks. Typical application that robots can realize include welding, painting, assembly, disassembly pick and place for printed circuit boards, packaging and labeling, palletizing, product inspection, and testing. Moreover, with the help of smart sensors, robots can ensure accomplish the tasks with high endurance, speed, and precision.
Cloud computing that brings computing services, including servers, storage, databases, softwar, etc. over the Internet, has become a economic and flexible computation resources deployment solution for enterprises, in recent years. Most cloud computing services fall into four broad categories: infrastructure as a service (IaaS), platform as a service (PaaS), serverless, and software as a service (SaaS) [43]. However, because cloud computing that resources are usually far away from the data sources, brings a long latency for IIoT data analysis, edge computing and fog computing that bring computation resources near data sources offer new ways of processing and analyzing data in real time. Currently, cloud-edge-fog architecture has become an effective architecture to provide sufficient computation and storage resources that are required by IIoT intelligence.

### 22 Feb 2024 v2

@@ Abstract
The fiercely competitive business environment and increasingly personalized customization needs are driving the digital transformation and upgrading of the manufacturing industry. IIoT intelligence, which can provide innovative and efficient solutions for various aspects of the manufacturing value chain, illuminates the path of transformation for the manufacturing industry. It’s time to provide a systematic vision of IIoT intelligence. However, existing surveys often focus on specific areas of IIoT intelligence, leading researchers and readers to have biases in their understanding of IIoT intelligence, that is, believing that research in one direction is the most important for the development of IIoT intelligence, while ignoring contributions from other directions. Therefore, this paper provides a comprehensive overview of IIoT intelligence. We first conduct an in-depth analysis of the inevitability of manufacturing transformation and study the successful experiences from the practices of Chinese enterprises. Then we give our definition of IIoT intelligence and demonstrate the value of IIoT intelligence for industries in fucntions, operations, deployments, and application. Afterwards, we propose a hierarchical development architecture for IIoT intelligence, which consists of five layers. The practical values of technical upgrades at each layer are illustrated by a close look on lighthouse factories. Following that, we identify seven kinds of technologies that accelerate the transformation of manufacturing, and clarify their contributions. The ethical implications and environmental impacts of adopting IIoT intelligence in manufacturing are analyzed as well. Finally, we explore the open challenges and development trends from four aspects to inspire future researches.
@@ I Introduction
Therefore, in this paper, we provide a comprehensive overview of IIoT intelligence. Considering that the development of IIoT intelligence integrates the wisdom of governments, enterprises, and researchers, we expand the information sources from academic papers to government documents, industry reports, and other related literature. Specifically, to extract insights into IIoT intelligence, we meticulously study the World Economic Forum’s Global Lighthouse Networks (GLN), which recognized the lighthouse use cases and factories under the Fourth Industrial Revolution. Then we try our best to carefully study associated customer stories from official sources. By delving into GLN and customer stories, we discern technologies pivotal for their digital and intelligent transformation. For each identified technology, we primarily examine technical papers published within last 5 years, as well as highly cited papers beyond this timeframe, to gather comprehensive insights and consolidate relevant knowledge. Finally, we present this paper to answer four questions as follows.
@@
To answer the second question, we propose a hierarchical development architecture to build a system-level understanding about how IIoT intelligence empowers smart manufacturing. Specifically, the architecture consists of five layers. Each layer has own missions. The equipment layer builds the basis of automatic production. The network layer connects human-machine-things, and overcomes isolated islands of information. The software layer provides digital representation of industrial processes. The modeling layer makes digital twin for physical processes to connect virtual and physical spaces. The analysis and optimization layer is responsible for mining industrial big data and optimizing industrial processes. Each layer of IIoT intelligence has its own key technologies, making its capabilities stronger. Interactions will occur between levels to fully enhance the intelligence and digitization level of the manufacturing industry. In addition, we also closely observe the lighthouse factory [28, 29, 30, 31] to demonstrate that technological upgrades at each layer enhance the company’s returns.
@@
This paper is structured as follows. Section II presents background and transformation secrets of smart manufacturing. The connotation of IIoT intelligence is explained in Section III. Section IV provides a reference hierarchical development architecture to build a systematic understanding for IIoT intelligence. Section V demonstrates the significant effects of IIoT intelligence through case studies on lighthouses and gradually released collaborations. Section VI identities some of the most influential technologies that facilitate IIoT intelligence. Section VII summarizes open challenges and future trends of IIoT intelligence. Section VIII presents the ethical implications and environmental impacts of adopting IIoT intelligence in manufacturing. The conclusion is given in Section IX.
@@ II-A Background
Recent years have witnessed the rapid development of a large number of advanced intelligent technologies, including cloud computing, 5G connection, industrial data analysis, industrial internet of things, etc. These technologies light the way for upgrading manufacturing. Many countries worldwide are also actively supporting the intelligent transformation of manufacturing through policy initiatives. These efforts encompass various aspects such as promoting research and development, fostering innovation ecosystems, incentivizing investment in advanced technologies, and addressing regulatory barriers. For example, on October 30 2023, the Biden Administration released Executive Order (E.O.) 14110 on Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence. It establishes a government-wide effort to guide responsible artificial intelligence (AI) development and deployment through federal agency leadership, regulation of industry, and engagement with international partners [33, 34]. On February 1st 2023, European commission released Green Deal Industrial Plan for the Net-Zero Age [35], and following with Net-Zero Industry Act on February 6 2024 [36], to help the European Union (EU) become home to clean technologies and make significant strides towards building a strong domestic manufacturing capacity of those technologies in the EU. On August 23, 2023, China released implementation guidelines as part of standards for new emerging industries, vowing to continuously improve the technical level and internationalization of new industry standards, and to provide solid technical support for accelerating the high-quality development of new industries by 2035 [37, 38].
@@ II-B Secrets to smart manufacturing
In this section, we will answer the question, how to accelerate the progress of intelligent transformation for manufacturing industry, based on the practical experiences of Foxconn Industrial Internet, which is a famous enterprise that has realized a high-level intelligent transformation in China. Please note that many other manufacturers, such as DHL Supply Chain, Siemens, and Unilever in America, Germany, and India, respectively, have undergone similar transformation journeys as Foxconn Industrial Internet. In this context, we draw insights from the experiences of Foxconn Industrial Internet to illustrate the secrets to smart manufacturing. Foxconn Industrial Internet has summarized its recipes into three pillars and two capabilities [39], as presented in Fig. 3.
@@ III-B How does IIoT intelligence achieve impacts in manufacturing?
IIoT intelligence can provide some innovative solutions for segments of the value chain of manufacturing, and support to build the pillars and capabilities (shown in Fig. 3). In the following sections, we delve into three case studies to illustrate the impacts of IIoT intelligence on smart manufacturing, focusing on economic, technical, and organizational aspects.
@@
III-B1 Economic impacts – reducing product R&D costs through digital experiments
@@
The traditional processes of the R&D of products require several rounds of trial production, testing and optimization in the physical production lines. Each round of the physical experiment will cost a lot of manpower, material, money and time. Therefore, the traditional R&D process is time-consuming, expensive and environmentally unfriendly.
IIoT intelligence can play a crucial role in reducing the costs of product R&D by digitizing a portion of physical experiments. Leveraging historical data and experiences, IIoT maps industrial processes into the digital realm, enabling cost-effective trial and error experiments. Moreover, the digital environment facilitates the automatic prediction of potential faults based on industrial knowledge and historical data. Recommendation technologies contribute by offering improvement suggestions, allowing the project team to promptly revise and optimize product designs in response to detected or predicted faults. This streamlined process enhances efficiency and cost-effectiveness in the development of new products.
In addition to optimizing the product development process and low-cost trial and error in the digital environment, IIoT intelligence can also support generative design in the product development stage, cross-departmental and cross-regional collaborative innovation and cooperation, thereby speeding up the design process and reducing design costs.
III-B2 Technical impacts – enhanced inspection efficiency through automated visual systems
Prior to the integration of intelligent technologies in manufacturing quality control, inspections relied on human inspectors for manual defect identification during visual inspections. However, this method had several drawbacks. Human inspectors may struggle to maintain consistent attention and precision over extended periods. Additionally, manual inspection processes may not match the speed of automated systems, resulting in slower production speeds and reduced efficiency.
IIoT intelligence tackles these challenges through automated visual inspection. Cameras and sensors strategically positioned along the production line capture images and gather data on product attributes. Machine vision systems, fueled by AI algorithms, process these images. Trained to recognize components, detect defects, and assess assembly quality, the algorithms prompt corrective actions when defects are found. The benefits of automated visual inspection include enhanced accuracy, a decrease in defect rates, and accelerated inspection processes, contributing to overall production line efficiency.
@@ III-B2 Technical impacts – enhanced inspection efficiency through automated visual systems
III-B3 Organizational behaviour impacts – increase labour efficiency through remote monitoring and control
Traditionally, workers are required to be physically present in the production area to actively engage in manufacturing processes or oversee equipment operations. This hands-on approach is essential for promptly identifying and rectifying errors during production. Additionally, equipment maintenance is typically carried out through on-site services provided by technicians. The deployment of production lines is often based on manual analysis and the expertise of industry professionals.
IIoT technologies can connect and perceive information across equipment, workers, and intermediate products in workshops. This enables remote monitoring of production conditions and control of manufacturing equipment, enhancing safety through unmanned operations and reducing exposure to hazardous sites. The technology improves Overall Equipment Effectiveness (OEE) and labor efficiency by transitioning from one-man-one-control to a more efficient one-man-multi-control approach. Remote monitoring and control allow managers to adapt worker schedules based on real-time progress. Additionally, these technologies create new white-collar job opportunities, enabling workers to operate from comfortable offices. This enhances the appeal of the manufacturing industry, addressing labor shortages by offering an alternative and attractive working environment.
@@ III-C The overall effect of IIoT intelligence on smart manufacturing
  • Functions: IIoT intelligence can help companies connect human-cyber-physical spaces, collect industrial data synchronously and accurately, build effective data analysis models, propose scenario-oriented optimization methods, develop microservice architecture, etc. For example, according to the lighthouse collections from the World Economic Forum [31] in 2023, through adopting IIoT intelligence, ACG Capsules in India realizes real-time quality batch insights and enhanced production planning and scheduling, which reduces 39% of batch lead time and improve 13% of on-time delivery in full.
  • Operation: IIoT intelligence builds cloud manufacturing business model for enterprises to enhance returns [41, 42, 43]. Cloud manufacturing create intelligent factory networks that encourage effective collaboration, through encouraging enterprises to encapsulate distributed manufacturing resources into cloud services and managing them in a centralized way. For example, China Resources Building Materials Technology adopts cloud to enrich and make data easy to consume across the organization to improve business performance. Agilent Technologies in Germany utilizes cloud to achieve predictive quality testing with AI, leading to a significant 13% improvement in test station throughput [31].
  • Deployment: IIoT intelligence provides the cloud-edge-terminal architecture to deploy computing resources in a cost-friendly way. Except that, IIoT intelligence can give efficient production deployment/scheduling suggestions to improve factory output and energy optimization. For example, ACG Capsules optimizes its production schedule with a novel color-matching AI and validates it with a digital twin, which realizes 10%-20% on-time-in-full increase.
  • Application scenarios: IIoT intelligence can be used to solve specific problems in some application scenarios, such as predictive and remote maintenance, product and equipment monitoring, product design feedback optimization, product process traceability, renting manufacturing capacity, product quality inspection, etc. For example, Johnson & Johnson Consumer Health in India deploys IIoT intelligence to achieve predictive maintenance, resulting in improved asset reliability and a 50% reduction in unplanned machine downtime [30].
@@ IV Hierarchical development architecture of IIoT intelligence
To build a systematic understanding of how IIoT intelligence deals with issues in the transformation process of smart manufacturing, this paper presents a hierarchical development architecture for IIoT intelligence. The architecture consists of equipment, network, software, model and optimization from bottom to top, as shown in Fig. 4. The classic architecture of IIoT-enabled smart manufacturing usually include the bottom two layers [9, 8, 44, 11], however, it cannot bear the weight of the various intelligence that have emerged in recent years. Therefore, we extends software layer, modeling layer, and analysis and optimization layer over the classic architecture. The functions of each layer in the hierarchical architecture are as follows:
@@
Compared with IIoT intelligence empowered smart manufacturing, traditional manufacturing models lack automatic controls and intelligent modeling. To be specific, traditional models often rely on manual control systems and fixed processes. Automation is limited, and adjustments to manufacturing processes typically require human intervention. This can lead to slower response times and increased susceptibility to human errors. When IIoT intelligence enabling smart manufacturing, IIoT sensors and devices are strategically deployed throughout the production process. These devices collect real-time data, enabling automatic control systems to respond dynamically to changing conditions. This results in enhanced automation, reduced manual intervention, and improved operational efficiency. Additionally, traditional models may lack sophisticated intelligent modeling capabilities. Decision-making often relies on historical data and experience, with limited ability to predict future events or proactively address issues. While IIoT intelligence provides advanced analytics and machine learning algorithms to process the vast amounts of data generated by sensors. These intelligent approaches help in forecasting maintenance needs, optimizing production schedules, and improving overall decision-making during the overall production processes.
@@ IV-A1 Motivations
Facing the crisis of labour shortage and high labour cost, the manufacturing industry’s desire for a machine that can work like or even better than humans has become increasingly strong, thus promoting the invention and innovation of industrial robots. Except that, there are other reasons that motivate enterprises [45]. a) There are many dull, dirty, dangerous and/or delicate tasks that can be done by robots, which can improve health, safety and job satisfaction of employees. b) The rapid development of robotics is continuously improving the automated and flexible production capacity of robots, while also reducing the procurement and maintenance costs. c) Modern robots will support a smaller carbon footprint [46][47], which is profitable for building low-carbon factories according to national requirements. d) More and more robots support plug-and-play integration and their control programming becomes much easy, which allow operators to become proficient in using robots with general training [48]. e) Human-robot collaboration will strengthen flexible production [49][50]. f) The aging society of the world has also released demands for liberation from manual labor.
Except robots, in order to achieve precise perception of the state of the production environment and high-precision control of the production process, it is necessary to introduce smart sensors to improve five senses111taste, touch, smell, sight, hearing of industry. In addition, with the deepening of industrial digital transformation, a large number of industrial data storage and analysis tasks have emerged. To complete these tasks, massive computing and storage resources are required.
@@ IV-A2 Connotation
An industrial robot refers to a robot system that can be used for manufacturing. Industrial robots are automated, programmable and capable of movement on three or more axes. Industrial robots are gradually replacing human to overtake dangerous and delicate tasks. Typical applications where robots can assist operations include welding, painting, assembly, disassembly pick and place for printed circuit boards, packaging and labeling, palletizing, product inspection, and testing. Moreover, with the help of smart sensors, robots can ensure accomplish tasks with high endurance, speed, and precision.
Cloud computing bringing computing services, including servers, storage, databases, softwar, etc. over the Internet, has become a economic and flexible computation resources deployment solution for enterprises in recent years. Most cloud computing services fall into four broad categories: infrastructure as a service (IaaS), platform as a service (PaaS), serverless, and software as a service (SaaS) [52]. However, because cloud resources are usually far away from the data sources, resulting in a long latency for IIoT data analysis, edge computing and fog computing that bring computation resources near data sources offer new ways of processing and analyzing data in real time. Currently, cloud-edge-fog architecture has become an effective architecture to provide sufficient computation and storage resources that are required by IIoT intelligence.